From GlueXWiki
Jump to: navigation, search

August 4, 2011 FDC meeting


  1. Production Construction Tracking (Dave)
    • Status
    • Wire deadening
    • Other
  2. Engineering update (Bill)
  3. Electronics update (Chris)
  4. Chamber testing at EEL126
    • Cell#1 status (Beni)
    • Testing prototype again (Lubomir)
  5. Other


Participants: Bill, Dave, Chris, Beni, Simon, Casey, and Lubomir.


- Dave: The construction tracking attached above has different structure now; it has the spare package with the first four wire frames (new epoxy used there) ready. Lubomir: as demonstrated with cell#2 just drying them for a week with nitrogen and/or Ar/CO2 would be enough to have them operational. Outside in the dirty area techs are laminating now the wire frames for the first package. After that (according to Dave) it will take less than a week per wire frame to finish the first package. Wire frame #5 was strung, wires taped and will be glued with the old Epoxy.

- Dave: we tested the electro-plating procedure this time on a full-scale wire frame, the one that was built at IUCF and shown on the open houses. Before that the frame was cleaned and HV applied to it in air. We were able to put on the two middle sectors 2200V with low current (~80 nA). After electroplating the current at 2200V in air almost didn't change. Two issues related to the procedure were discussed:

  • First, how to bypass the 1MOhm resistors on each wire and the 10KOhm on the HV sector; this is needed to use the HV paths to connect to the sense wires. Since we are plating only 6 or 8 wires this is not a big deal. Lubomir: the problem is that if by some reasons you had a bad connection to a wire and you don't know what was the current there, you can't repeat the electroplating to fix that wire. One way would be to unsolder the resistors and shorten the pads (as Dave did in this test), or simply to solder a wire on the top of the resistors, or to use a small alligator clamp (Chris) across the resistor. Bill proposed another option: after stringing not to cut these wires at the soldering pad, but to leave some length and use it to apply the voltage there. Related to that, Beni figured out that actually the current trough each individual wire is different, proportional the wire length that is in the chemical and that's why we have the same thickening on all the wires (Dave confirmed it by measurements with micrometer on the small prototype frames).
  • Second, how to check the results of the wire deadening. So far, with the full-scale frame, we were not able to look at the wires with good enough magnification. Using microscope in the middle of the wires is not practical. We will try to use the position measurement system to see if the laser sensor can measure the thickness of the wires, at least relatively.

- Lubomir: wire frame #2 was keeping the HV (+2200/-500V) without any trips for almost a week (except some intentional interruptions) and the current is relatively low, ~200nA for the whole chamber. Before that it took ~5 days drying with nitrogen and the current between the field and sense wires was going down, but after changing the gas to Ar/CO2 the field-sense wire current dropped significantly from ~250nA to 20nA. Most likely there was some moister in the nitrogen bottle.


- Bill: we received the strong-back for the package; ordered legs for it (ready in two days). Also ordered second set of gusset rings and hubs needed for the spare package, to be ready in two weeks. Simon: it will be good to put the strong-back on the chamber in 126 to see if it will solve the flatness problem.


- Chris: found again problems on several connectors on the daughter cards: solder bridges between the pins. Will inform the company, but actually all the cards are produced already and according to Chris it is easy to fix them. We have now all the signal cables, testing in the moment FDC type 4. Storage space for the cables was found in the ARC building.

Chamber testing

- Beni: DAQ is now working with the modified VXS back plane. More studies showed that the drift time spectra looks very different from what we had with the prototype for 90/10 Ar/CO2. Most likely in the first chamber we have problem with the gas mixture in addition to the flatness and HV problems. Also Beni found that the efficiency of the chamber now is ~60% and was 99% before, but it could be a result of the new DAQ and/or related software. For that, it would be good to have a working chamber to troubleshoot these problems.

- Beni on testing Humi-seal with radiation: the effect we expect to see is the polymerization on the wires (Molter effect) and it happens when both: there's high radiation, and the HV is on for a long period of time. So, if we want to see these effects it will take long time and high radiation doses. However, it is known that one can use alcohol in the gas mixture to prolong the chamber life. Therefore, we should start these tests but at the same time we can use Humi-seal in the chambers production and, if we have to, we will add alcohol to the gas mixture.

- After the meeting there were additional discussions. It is clear that we want to have a working production chamber at 126 and at the same time to do these long-term studies with Humi-seal. Going back to the prototype to study only the Humi-seal would not be an efficient solution. That's why it was proposed to replace the wire frame #1 with wire frame #2 and test it at 126. Frame #2 also has Humi-seal on the capacitors and we can do all the tests simultaneously. In addition we can add the strong-back to that package to test the flatness.